# Artificial Transmission Line Synthesis Tailored for Traveling-Wave Parametric Processes

### v1

@@ Abstract
Artificial transmission lines built with lumped-element inductors and capacitors form the backbone of broadband, nearly quantum-limited traveling-wave parametric amplifiers (TWPAs). However, systematic design methods for TWPAs, and more generally artificial transmission lines, are lacking. Here, I develop a general synthesis framework for lossless artificial transmission lines by borrowing from periodic structure theory and passive network synthesis. These complementary approaches divide the design space: periodic loading synthesis employs spatial modulation of frequency-independent components, while filter synthesis employs frequency-dependent responses in spatially-uniform components. When tailoring transmission lines for parametric processes, nonlinear elements are added, typically nonlinear inductances in superconducting circuits, while ensuring energy and momentum conservation between interacting tones. Applying this framework, I design a kinetic inductance TWPA with a novel phase-matching architecture, and a backward-pumped Josephson TWPA exploiting an ambidextrous i.e., right-left-handed transmission line.
@@ I Introduction
In linear circuit design, synthesis and analysis are complementary. Analysis computes the response of a given network, for example by cascading ABCD matrices [1], whereas synthesis tackles the inverse problem: systematically designing a network that approximates a desired response [2]. Though synthesis does not guarantee the resulting network is unique or optimal, it provides a systematic framework that reveals what responses are achievable without trial and error. Beyond design efficiency, synthesis can unveil previously unknown network topologies, enabling the exploration of exotic device concepts that emerge naturally from the framework.
While resonant parametric amplifiers have recently benefited from systematic synthesis approaches for gain and bandwidth [3], traveling-wave parametric amplifiers (TWPAs), essential components in superconducting quantum circuit readout [4, 5, 6, 7, 8], lack equivalent design tools. This absence stems partly from the lack of systematic synthesis methods for the artificial transmission lines (ATLs) that form the backbone of TWPA architectures.
Having a synthesis framework would enable a more systematic design of TWPA bandwidths and dispersion relations. When designing a TWPA, a key challenge is to selectively enable mode propagation in order to favor desired parametric processes, while suppressing spurious processes that deplete pump photons and degrade noise performance [9, 10]. A phase-mismatch will prevent the exponential growth of a given process along the ATL, but to eliminate it completely one of the modes involved in the process must not propagate. Mode filtering becomes all the more critical in devices leveraging several parametric processes, for example parametric amplification and frequency conversion [11, 12], where multiple pump tones can mix together. Mitigation strategies exist, ranging from leveraging natural chromatic dispersion [13, 14, 15], to periodic loading techniques [16, 17, 18, 19, 20, 21] and Floquet mode propagation [10, 22], but they are generally engineered by trial and error, iterating on circuit component values to obtain a desired response.
Here, I develop a comprehensive synthesis approach for lossless ATLs by recognizing that the design space is naturally divided into two complementary approaches: periodic loading, where frequency-independent components vary spatially, and filter synthesis, where spatially-invariant components exhibit frequency-dependent responses. In Sec. II, I present this theoretical framework, developing both approaches and their fundamental constraints. In Sec. III, I apply these methods to design two TWPAs: a kinetic inductance device with a novel phase-matching structure, and a backward-pumped Josephson TWPA exploiting “ambidextrous” transmission line behavior. I conclude by discussing the broader implications of systematic ATL synthesis for superconducting quantum circuits.
@@ II.1 General form for an artificial transmission line
In the context of electrical circuits, an ATL is a one-dimensional periodic lattice built with lumped circuit elements, supporting the propagation of voltage and current waves. Topologically, a lossless ATL consists of an infinite cascade of series reactances, typically inductors, alternating with susceptances, typically capacitors, shunted to ground. A unit cell within the ATL is composed of an inductor-capacitor pair; it is antimetric when the entire inductor precedes or succeeds the capacitor, and symmetric when either the inductor or the capacitor is halved and surrounds the other element within the pair, see Fig. 1a. Both inductors and capacitors can have a periodic spatial dependency. Both elements can also be replaced with sub-networks of inductors and capacitors. This leads to two ways of designing the response of an ATL, see Fig. 1b: engineer the periodic modulation of inductors and capacitors, or engineer the frequency response of spatially-invariant reactances and susceptances. Below, I present a synthesis framework for both approaches separately, and then discuss the effect of combining them within the same lattice.
Figure 1: General topology of an artificial transmission line. (a) Cascaded unit cells, comprising series reactances (typically inductors) and shunt susceptances (typically capacitors), can be made antimetric (blue) or symmetric (orange). (b) The synthesis framework divides into two approaches: periodic loading (left) where the inductors and capacitors vary spatially, and filter synthesis (right) where the spatially-invariant reactances (X) and susceptances (B) are engineered to give a specific frequency response.
@@ II.2 Periodic loading synthesis
Spatially modulating frequency-independent components creates stopbands in the ATL response whose systematic design requires a generalized dispersion relation that captures the effect of periodic loading. I derive this relation in Sec. II.2.1, then show in Sec. II.2.2 how to synthesize stopbands by specifying their center frequencies and widths.
@@ II.2.1 Dispersion relation
where $k_{d}=2\pi/d$. Note that since $L$ and $C$ are real, $l_{n}=l_{-n}$ and $c_{n}=c_{-n}$ for $n\in\mathbb{Z}$.
@@
which is the generalized dispersion relation of a periodically loaded ATL. Here, $\mathbf{M}=\tilde{\omega}^{2}\mathbf{L}\mathbf{C}-\mathbf{K}^{2}$ defines the coupling matrix between the spatial harmonics $\{e^{-jnk_{d}x}\}_{n\in\mathbb{Z}}$.
@@ II.2.2 Stopband synthesis
Figure 2: Example of stopband synthesis via periodic modulation of the ATL shunt capacitance. (a) The band diagram shows the dispersion relation for the periodically loaded ATL with two synthesized stopbands at $8$ and $24$ GHz. Each line represents a different spatial harmonic $n$, with blue lines indicating modes where $k+nk_{d}>0$ (forward propagation) and orange lines where $k+nk_{d}<0$ (backward propagation). The stopbands appear at the edge of the Brillouin zones where counter-propagating modes couple, creating the bandgaps shown in gray. The inset shows the normalized capacitance modulation profile over one supercell ($d=59$ cells), consisting of two overlapped cosine modulations. (b) Zoom of the first stopband at $8$ GHz with width $\Delta_{1}^{+}=0.2$ GHz. (c) Zoom of the third stopband at $24$ GHz with width $\Delta_{3}^{-}=2$ GHz.
@@
$\begin{vmatrix}\vdots&\vdots&\ddots&\vdots&\vdots\\ \cdots\!&\tilde{\omega}^{2}-(\tilde{k}-1)^{2}&\tilde{\omega}^{2}c_{1}&\tilde{\omega}^{2}c_{2}&\cdots\!\\ \cdots\!&\tilde{\omega}^{2}c_{1}&\tilde{\omega}^{2}-\tilde{k}^{2}&\tilde{\omega}^{2}c_{1}&\cdots\!\\ \cdots\!&\tilde{\omega}^{2}c_{2}&\tilde{\omega}^{2}c_{1}&\tilde{\omega}^{2}-(\tilde{k}+1)^{2}&\cdots\!\\ \vdots&\vdots&\ddots&\vdots&\vdots\end{vmatrix}$
@@
Even though $\mathbf{M}$ is an infinite matrix, it can be truncated because the stopband structure is fully determined by the coupling between harmonics in the range $n\in[-N,N]$, where $N$ corresponds to the highest stopband frequency. When designing several stopbands, $\tilde{k}_{1}=1/2$ is associated to the greatest common divisor (gcd) frequency $\tilde{\omega}_{1}$ of all the stopband frequencies. In other words, $d$ is chosen so that all the stopband frequencies can be decomposed onto the same spatial harmonic basis. In that case, $\mathbf{M}$ may be sparse, because only the harmonics corresponding to a stopband create non-zero off-diagaonal terms.
@@
[…]
@@ II.3.2 Filter transformations: engineering any pass-stop response
$k_{2i}$ $=\left.(s+j\omega_{2i})\lambda(s)\right\rvert_{s=j\omega_{2i}}$ (22)
@@ II.4.2 Practical implementations
In practice, the phase-matching condition in TWPAs can be achieved via periodic loading [16, 17, 18, 29, 21], but it presents two issues: (i) in a finite ATL, it creates ripples near the stopband [19], where the optimal pump frequency for phase-matching usually is. Thus, the phase-matched pump may suffer from an unfavorable impedance mismatch, which leads to pump reflections and backward gain. Furthermore, these ripples may land within the amplification band, affecting the gain profile. Apodization techniques can mitigate these spurious effects [29]. (ii) When generated from periodic loading, the phase shift near a stopband only slowly approaches the edge of the first Brillouin zone $\pm\pi/d$, with $d$ the number of cells per supercell. This phase shift is limited by the loading value. In contrast, filter synthesis can achieve a sharp jump to the same $\pm\pi/d$ limit, but the transition sharpness is controlled by the pole-zero separation rather than spatial harmonic coupling (see Fig. 5c). Consequently, achieving phase-matching via periodic loading requires heavily loading the ATL, creating large stopband regions unusable for amplification, while filter synthesis provides localized phase shift, with minimal bandwidth penalty.
@@ III.2.1 Backward phase-matching
In a 3WM TWPA the phase-matching condition follows directly from the linear dispersion relation: $k_{p}=k_{s}+k_{i}$, where $k_{p}$, $k_{s}$, and $k_{i}$ are the wavenumbers of the pump, signal and idler frequencies, respectively. Unlike 4WM, which requires nonlinear corrections, it allows direct use of the band diagram from the ATL synthesis framework to directly identify where and how this wavenumber relationship can be satisfied. Additionally, the pump frequency naturally falls above the amplification band, eliminating pump filtering concerns.
@@ IV Conclusion
I have developed a systematic framework encompassing two complementary approaches for designing artificial transmission lines: periodic loading synthesis, where frequency-independent components are spatially modulated to create stopbands in right-handed ATLs, and filter synthesis, where spatially-invariant components are engineered with frequency-dependent responses. Periodic loading synthesis relies on decomposing voltage, current, and component modulations onto a common spatial harmonic basis, which reveals a generalized matrix dispersion relation that can be used for systematic stopband design. ATL synthesis builds on filter theory: recognizing that a right-handed ATL behaves as a low-pass prototype enables direct application of filter transformations to ATL design.
This unified synthesis framework reveals fundamental constraints on ATL dispersion relations: the wavenumber $k$ is always a monotonically increasing function of frequency, though not necessarily continuous, a result that parallels Foster’s reactance theorem in the context of artificial transmission lines. This theoretical foundation enables systematic design, revealing previously unexplored parametric amplifier architectures that emerge naturally from the synthesis procedure.
I validated this approach by designing and simulating two novel TWPAs using an extended open-source harmonic balance solver that supports arbitrary nonlinear inductances defined through their inductance-flux Taylor expansion. The first, a 4WM KTWPA, employs a new rpm filter topology particularly suited for kinetic inductance devices. The second demonstrates a new concept: a backward-pumped TWPA (b-TWPA) based on an ambidextrous transmission line where the pump propagates backward while signal and idler propagate forward. More broadly, such ambidextrous ATLs enable exotic phase-matching schemes, including configurations that could generate entangled counter-propagating photon pairs—a phenomenon well-established in optics but unexplored in superconducting microwave circuits.
The significance of systematic synthesis extends beyond design efficiency. By revealing what dispersion behaviors are allowed, the framework uncovers new physical phenomena and device concepts that might otherwise remain hidden. Superconducting microwave circuits offer unique advantages for exploring these concepts: they can be constructed from the ground up using lumped inductors and capacitors while maintaining ultra-low loss. This opens possibilities for applications ranging from quantum-limited amplification to exotic photon state generation.

### v2

@@ Abstract
Artificial transmission lines built with lumped-element inductors and capacitors form the backbone of broadband, nearly quantum-limited traveling-wave parametric amplifiers (TWPAs). When tailoring these transmission lines for parametric processes, nonlinear elements are added, typically nonlinear inductances in superconducting circuits, and energy and momentum conservation between interacting tones must be enforced through careful design of the ATL dispersion relation. However, a unified theoretical framework describing achievable dispersion relations is lacking. Here, I develop such a framework, borrowing from periodic structure theory and passive network synthesis. These complementary approaches divide the design space: periodic loading synthesis employs spatial modulation of frequency-independent components, while filter synthesis employs frequency-dependent responses in spatially-uniform components. The framework reveals fundamental constraints and enables the discovery of novel TWPA architectures. In particular, I design a kinetic inductance TWPA with a novel phase-matching architecture, and a backward-pumped Josephson TWPA exploiting an ambidextrous i.e., right-left-handed transmission line.
@@ I Introduction
In linear circuit design, synthesis and analysis are complementary. Analysis computes the response of a given network, for example by cascading ABCD matrices [1], whereas synthesis tackles the inverse problem: designing a network that approximates a desired response [2]. Though synthesis does not guarantee the resulting network is unique or optimal, it is built on a theoretical framework that reveals what responses are achievable and what network topologies are possible, enabling the exploration of exotic device concepts.
While resonant parametric amplifiers have recently benefited from systematic synthesis approaches for gain and bandwidth [3], traveling-wave parametric amplifiers (TWPAs), essential components in superconducting quantum circuit readout [4, 5, 6, 7, 8], lack equivalent design tools. This absence stems partly from the lack of such a framework describing the dispersion of artificial transmission lines (ATLs) that form the backbone of TWPA architectures.
Controlling the ATL’s dispersion relation is essential when designing a TWPA, because a key challenge is to selectively enable mode propagation in order to favor desired parametric processes, while suppressing spurious processes that deplete pump photons and degrade noise performance [9, 10]. A phase-mismatch will prevent the exponential growth of a given process along the ATL, but to effectively suppress it, a stopband can be engineered to prevent the coherent buildup of one of the modes involved. Mode filtering becomes all the more critical in devices leveraging several parametric processes, for example parametric amplification and frequency conversion [11, 12], where multiple pump tones can mix together and produce larger sets of spurious processes. Heuristic mitigation strategies exist, ranging from leveraging natural chromatic dispersion [13, 14, 15], to periodic loading techniques [16, 17, 18, 19, 20, 21] and Floquet mode propagation [10, 22].
Here, I develop a theoretical framework for lossless ATLs by recognizing that the design space is divided into two complementary synthesis approaches: periodic loading, where frequency-independent components vary spatially, and filter synthesis, where spatially-invariant components exhibit frequency-dependent responses. In Sec. II, I present this theoretical framework, developing both approaches and their fundamental constraints. In Sec. III, I apply these methods to design two TWPAs: a kinetic inductance device with a novel phase-matching structure, and a backward-pumped Josephson TWPA exploiting “ambidextrous” transmission line behavior. I conclude by discussing the broader implications of this framework for superconducting quantum circuits.
@@ II.1 General form for an artificial transmission line
In the context of electrical circuits, an ATL is a one-dimensional periodic lattice built with lumped circuit elements, supporting the propagation of voltage and current waves. Topologically, a lossless, right-handed ATL consists of an infinite cascade of series reactances, typically inductors, alternating with susceptances, typically capacitors, shunted to ground. A unit cell within the ATL is composed of an inductor-capacitor pair; it is antimetric when the entire inductor precedes or succeeds the capacitor, and symmetric when either the inductor or the capacitor is halved and surrounds the other element within the pair, see Fig. 1a. Both inductors and capacitors can have a periodic spatial dependency. Both elements can also be replaced with sub-networks of inductors and capacitors. This leads to two ways of designing the response of an ATL, see Fig. 1b: engineer the periodic modulation of inductors and capacitors, or engineer the frequency response of spatially-invariant reactances and susceptances. Below, I present a synthesis framework for both approaches separately, and then discuss the effect of combining them within the same lattice.
Figure 1: General topology of a right-handed artificial transmission line. (a) Cascaded unit cells, comprising series reactances (typically inductors) and shunt susceptances (typically capacitors), can be made antimetric (blue) or symmetric (orange). (b) The synthesis framework divides into two approaches: periodic loading (left) where the inductors and capacitors vary spatially, and filter synthesis (right) where the spatially-invariant reactances (X) and susceptances (B) are engineered to give a specific frequency response.
@@ II.2 Periodic loading synthesis
Spatially modulating frequency-independent components creates stopbands in the ATL response whose systematic design requires a generalized dispersion relation that captures the effect of periodic loading. I derive this relation in Sec. II.2.1, then show in Sec. II.2.2 how to synthesize stopbands by specifying their center frequencies and widths. Here, the right-handed ATL can be viewed as a prototype; other topologies, including left-handed and composite lines, can be obtained by applying the frequency transformations of Sec. II.3.2 to this prototype (see Sec. II.4.1).
@@ II.2.1 Dispersion relation
where $k_{d}=2\pi/d$. Note that since $L$ and $C$ are real-valued functions, $l_{n}=l_{-n}^{*}$ and $c_{n}=c_{-n}^{*}$ for $n\in\mathbb{Z}$.
@@
which is the generalized dispersion relation of a periodically loaded, right-handed ATL. Here, $\mathbf{M}=\tilde{\omega}^{2}\mathbf{L}\mathbf{C}-\mathbf{K}^{2}$ defines the coupling matrix between the spatial harmonics $\{e^{-jnk_{d}x}\}_{n\in\mathbb{Z}}$.
@@ II.2.2 Stopband synthesis
Figure 2: Example of stopband synthesis via periodic modulation of the ATL shunt capacitance. (a) The band diagram shows the dispersion relation for the periodically loaded ATL with two synthesized stopbands at $8$ and $24$ GHz. Each line represents a different spatial harmonic $n$, with blue lines indicating modes where $k+nk_{d}>0$ (forward propagation) and orange lines where $k+nk_{d}<0$ (backward propagation). The stopbands appear at the edge of the Brillouin zones where counter-propagating modes couple, creating the bandgaps shown in gray. The inset shows the capacitance modulation profile over one supercell ($d=59$ cells), consisting of two overlapped cosine modulations. (b) Zoom of the first stopband at $8$ GHz with width $\Delta_{1}^{+}=0.2$ GHz. (c) Zoom of the third stopband at $24$ GHz with width $\Delta_{3}^{-}=2$ GHz.
@@
$\begin{vmatrix}\vdots&\vdots&\ddots&\vdots&\vdots\\ \cdots\!&\tilde{\omega}^{2}-(\tilde{k}-1)^{2}&\tilde{\omega}^{2}c_{1}&\tilde{\omega}^{2}c_{2}&\cdots\!\\ \cdots\!&\tilde{\omega}^{2}c_{1}^{*}&\tilde{\omega}^{2}-\tilde{k}^{2}&\tilde{\omega}^{2}c_{1}&\cdots\!\\ \cdots\!&\tilde{\omega}^{2}c_{2}^{*}&\tilde{\omega}^{2}c_{1}^{*}&\tilde{\omega}^{2}-(\tilde{k}+1)^{2}&\cdots\!\\ \vdots&\vdots&\ddots&\vdots&\vdots\end{vmatrix}$
@@
Even though $\mathbf{M}$ is an infinite matrix, it can be truncated because the stopband structure is fully determined by the coupling between harmonics in the range $n\in[-N,N]$, where $N$ corresponds to the highest stopband frequency. When designing several stopbands, $\tilde{k}_{1}=1/2$ is associated to the greatest common divisor (gcd) frequency $\tilde{\omega}_{1}$ of all the stopband frequencies. In other words, the number $d$ of cells per supercell is chosen so that all the stopband frequencies can be decomposed onto the same spatial harmonic basis. In that case, $\mathbf{M}$ may be sparse, because only the harmonics corresponding to a stopband create non-zero off-diagonal terms.
@@
The constraints generating $\{D_{n}=0\}_{n\in N}$ may not always yield realizable solutions. In fact, physical realizability requires $\lvert c_{n}\rvert\leq 0.5$ (necessary but not generally sufficient), because $C$ cannot take negative values (since $C=C_{0}[1+c_{1}e^{-jk_{d}x}+c_{1}^{*}e^{jk_{d}x}+\cdots$], see Eq. 2). In other words, the stopbands have a maximum width. For the case of a single stopband at $\tilde{k}=1/2$, where only the $n=0$ and n=1n=-1 harmonics couple, the relative stopband width $\delta_{1}=(\Delta_{1}^{+}+\Delta_{1}^{-})/f_{1}$ can be derived analytically; $\delta_{1}^{\mathrm{max}}\simeq 0.6$ when $c_{1}=0.5$, see appendix B.
@@ II.3.2 Filter transformations: engineering any pass-stop response
$k_{2i}$ $=\left.(s-j\omega_{2i})\lambda(s)\right\rvert_{s=j\omega_{2i}}$ (22)
@@ II.4.2 Practical implementations
In practice, the phase-matching condition in TWPAs can be achieved via periodic loading [16, 17, 18, 29, 21], but it presents two issues: (i) in a finite ATL, it creates ripples near the stopband [19], where the optimal pump frequency for phase-matching usually is. Thus, the phase-matched pump may suffer from an unfavorable impedance mismatch, which leads to pump reflections and backward gain. Furthermore, these ripples may land within the amplification band, affecting the gain profile. Apodization techniques can mitigate these spurious effects [29]. (ii) When generated from periodic loading, the phase shift near a stopband only slowly approaches the edge of the first Brillouin zone $\pm\pi/d$, with $d$ the number of cells per supercell. This phase shift is limited by the loading value. In contrast, filter synthesis can achieve a sharp jump to the same $\pm\pi/d$ limit, but the transition sharpness is controlled by the pole-zero separation, which in practice is constrained by achievable component values, rather than spatial harmonic coupling (see Fig. 5c). Consequently, achieving phase-matching via periodic loading requires heavily loading the ATL, creating large stopband regions unusable for amplification, while filter synthesis provides localized phase shift, with a bandwidth penalty set by the pole-zero separation.
@@ III.2.1 Backward phase-matching
In a 3WM TWPA the phase-matching condition follows directly from the linear dispersion relation: $k_{p}=k_{s}+k_{i}$, where $k_{p}$, $k_{s}$, and $k_{i}$ are the wavenumbers of the pump, signal and idler frequencies, respectively. Unlike 4WM, which requires nonlinear corrections, it allows direct use of the band diagram from the ATL synthesis framework to directly identify where and how this wavenumber relationship can be satisfied. Additionally, the pump frequency naturally falls above the amplification band, significantly relaxing pump filtering requirements.
@@ IV Conclusion
I have developed a theoretical framework for designing artificial transmission lines, encompassing two complementary synthesis approaches: periodic loading synthesis, where frequency-independent components are spatially modulated to create stopbands in right-handed ATLs, and filter synthesis, where spatially-invariant components are engineered with frequency-dependent responses. Periodic loading synthesis relies on a generalized matrix dispersion relation that can be used for stopband design, found by decomposing voltage, current, and component modulations onto a common spatial harmonic basis. ATL design via filter synthesis recognizes that a right-handed ATL behaves as a low-pass prototype, enabling direct application of filter transformations.
This framework reveals fundamental constraints on ATL dispersion relations: in particular, the wavenumber $k$ is always a monotonically increasing function of frequency, though not necessarily continuous, a result that parallels Foster’s reactance theorem in the context of artificial transmission lines. It also reveals previously unexplored parametric amplifier architectures.
I validated this approach by designing and simulating two novel TWPAs using an extended open-source harmonic balance solver that supports arbitrary nonlinear inductances defined through their inductance-flux Taylor expansion. The first, a 4WM KTWPA, employs a new rpm filter topology particularly suited for kinetic inductance devices. The second demonstrates a new concept: a backward-pumped TWPA (b-TWPA) based on an ambidextrous transmission line where the pump propagates backward while signal and idler propagate forward.
Superconducting circuits offer unique advantages for exploring these new concepts: they can be constructed from the ground up using lumped inductors and capacitors while maintaining ultra-low loss. This flexibility, however, generates a vast parameter space. A theoretical framework that constrains this space, such as the one presented here, is therefore a valuable tool for navigating the design process.
@@ Appendix I Component values for TWPA designs
Appendix J Third-harmonic suppression in the 4WM KTWPA
Figure 11 compares two variants of the KTWPA presented in Sec. III.1, with the rpm features removed to isolate the effect of the third-harmonic stopband: one retaining the stopband at $3f_{p}$ and one without. Without the stopband, the 2nd and 3rd pump harmonics reach peak powers of 67-67 dBm and 78-78 dBm respectively, compared to 88-88 dBm and 105-105 dBm with the stopband, a suppression of approximately 20–27 dB. Both designs exhibit similar pump power variations ($\sim 28\,\%$) along the line, indicating that the stopband does not primarily prevent pump depletion. Rather, its essential role is to enforce that amplification occurs exclusively through the designed 4WM process: without the stopband, the signal gain profile becomes strongly asymmetric about the pump frequency, and the quantum efficiency degrades from >0.95>0.95 to 0.520.520.760.76.
Figure 11: Effect of the third-harmonic stopband on the KTWPA of Sec. III.1, with the rpm features removed. The left column (a,c,e) corresponds to the design with the stopband at $3f_{p}$, and the right column (b,d,f) to the design without. (a,b) Spatial power distribution of the pump fundamental ($f_{p}$, purple), second harmonic ($2f_{p}$, orange), and third harmonic ($3f_{p}$, pink) along the device. Without the stopband, the 2nd and 3rd harmonics reach peak powers 20–27 dB higher. (c,d) Signal gain $|S_{21}|^{2}$. With the stopband, the gain profile is symmetric about the pump frequency, characteristic of the 4WM process. Without it, the profile becomes strongly asymmetric, indicating that parasitic higher-order parametric processes dominate the amplification. (e,f) Quantum efficiency normalized to the theoretical maximum. With the stopband, $\mathrm{QE}/\mathrm{QE_{ideal}}$ remains above $0.95$ across the signal band; without it, it degrades to 0.520.520.760.76.